Simplify. Please show all of your work.
-w + 12w2 – 15w3 + 3(12w + 3w2 – 5w3)

Answer by ruru7723(1)

-w+12w2-15w3+3(12w+3w2-5w3)=
-w+12w2-15w3+36w+9w2-15w3=
35w+21w2